A non-relativistic free electron has kinetic energy K. If its wavelength λ triples, its kinetic energy is:


A.
K/4

B.
6 K

C.
K/3

D.
K/9

E.
3 K
1
Expert's answer
2018-04-09T03:15:14-0400
K=p^2/2m
p=hc/λ~1/λ
Thus,
K~1/λ^2
K^'/K=1/3^2 =1/9

Answer D. K/9